Limit x approach infinity x[ ( 1 + a/x )raise to power of 1+1/x - x raise to power of -1/x ( x+ a ) ]



1
Expert's answer
2021-01-18T06:57:31-0500

Here we have to solve "\lim\limits_{x\to\infin}(1+\frac{a}{x})^x"


Let us assume

"y=(1+\frac{a}{x})^x"

Taking natural logarithm on both sides we get,



"\Rightarrow\ln(y)=\ln[(1+\frac{a}{x})^x]\\
\Rightarrow \ln(y)=x\ln(1+\frac{a}{x})"

Now, taking "\lim\limits_{x\to\infin}" on both sides we have,



"\Rightarrow \lim\limits_{x\to\infin}\ln(y)=\lim\limits_{x\to\infin}x\ln(1+\frac{a}{x})\\""\Rightarrow\lim\limits_{x\to\infin}\ln(y)=\lim\limits_{x\to\infin}\frac{ln(1+\frac{a}{x})}{\frac{1}{x}}"

Using L'Hospital's rule to solve the RHS of the equation, as the limit is in the form of "\frac{0}{0}" we have,



"\Rightarrow\lim\limits_{x\to\infin}\ln(y)=\lim\limits_{x\to\infin}\frac{\frac{1}{1+\frac{a}{x}}}{\frac{-1}{x^2}}(\frac{-a}{x^2})"

"\Rightarrow\lim\limits_{x\to\infin}\ln(y)=a\\"

Now, as we can interchange "\ln" and "\lim" we get,



"\Rightarrow\ln\lim\limits_{x\to\infin}(y)=a\\
\Rightarrow\lim\limits_{x\to\infin}y=e^a\\
\Rightarrow\lim\limits_{x\to\infin}(1+\frac{a}{x})^x=e^a"